OpenStudy (calculusxy):

MEDAL!!! I need help with graphing, putting from standard to vertex form, and writing its transformation. \(y = x^2 - 8x + 12\)

OpenStudy (calculusxy):

@SolomonZelman

OpenStudy (calculusxy):

i know vertex form is: \(a(x - h)^2 + k\) h = -b/2a = 8/2 = 4 k = 16 - 32 + 12 k = -4 vertex form is \((x - 4)^2 - 4\)?

OpenStudy (civicsiscool44):

yes whats your Q

OpenStudy (civicsiscool44):

you want answer?

OpenStudy (calculusxy):

no i want help on getting to the answer

OpenStudy (civicsiscool44):

ok

OpenStudy (calculusxy):

i really need help on transformation.

OpenStudy (civicsiscool44):

sorry i can't help you there

OpenStudy (calculusxy):

ok but can u make sure my vertex form?

OpenStudy (civicsiscool44):

it looks correct

OpenStudy (calculusxy):

thanks!

OpenStudy (calculusxy):

@SolomonZelman

OpenStudy (civicsiscool44):

@BeautifulMystery.